### v1

# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

Abstract.
Wi-Fi sensing has emerged as a transformative technology that leverages ubiquitous wireless signals to enable a variety of applications ranging from activity and gesture recognition to indoor localization and health monitoring. However, the inherent dependency of Wi-Fi signals on environmental conditions introduces significant generalization challenges—variations in surroundings, human positions, and orientations often lead to inconsistent signal features, impeding robust action recognition.
In this survey, we review over 200 studies on Wi-Fi sensing generalization, categorizing them along the entire sensing pipeline: device deployment, signal processing, feature learning, and model deployment. We systematically analyze state-of-the-art techniques, which are employed to mitigate the adverse effects of environmental variability. Moreover, we provide a comprehensive overview of open-source datasets such as Widar3.0, XRF55, and XRFv2, highlighting their unique characteristics and applicability for multimodal fusion and cross-modal tasks. Finally, we discuss emerging research directions—including multimodal approaches and the integration of large language models—to inspire future advancements in this rapidly evolving field. Our survey aims to serve as a valuable resource for researchers, offering insights into current methodologies, available datasets, and promising avenues for further investigation.
Wi-Fi sensing, Human sensing, Action recognition, Indoor localization, Domain adaptation, Domain generalization, Few-shot learning, Generative adversarial network, Transfer learning, Meta-learning
copyright: acmlicensedjournal: CSURjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviews
1. Introduction
Wi-Fi sensing has emerged as a transformative technology that exploits the pervasive Wi-Fi infrastructure for purposes beyond traditional communication. By examining fluctuations in Wi-Fi signals, particularly Channel State Information (CSI) and Received Signal Strength Indicator (RSSI), researchers have explored its potential across diverse applications, including activity recognition (Jiang et al., 2018b; Li et al., 2021d; Wang et al., 2017a, 2015; Tan et al., 2019; Zhu et al., 2017b; Wang et al., 2014; Chen et al., 2018; Adib and Katabi, 2013; Qian et al., 2017b), gesture recognition (Li et al., 2016b; Pu et al., 2013; Tan and Yang, 2016; Li et al., 2016c; Ali et al., 2015; He et al., 2015; Venkatnarayan et al., 2018; Abdelnasser et al., 2015; Ma et al., 2018; Wang et al., 2019a; Sun et al., 2015), indoor localization (Li et al., 2016a; Xie et al., 2015; Wang et al., 2016a; Vasisht et al., 2016; Qian et al., 2017a; Gjengset et al., 2014; Li et al., 2017; Xie et al., 2019; Qian et al., 2018; Kotaru et al., 2015), user authentication and identification (Korany et al., 2019; Shahzad and Zhang, 2018; Wang et al., 2019c, b; Shi et al., 2017; Zeng et al., 2016; Zhang et al., 2020c; Wang et al., 2016b), health monitoring (Zeng et al., 2020; Wang et al., 2017b; Zeng et al., 2018; Wang et al., 2016d; Zheng et al., 2016; Liu et al., 2014; Zeng et al., 2019), fallen detection (Wang et al., 2016f; Palipana et al., 2018; Wang et al., 2016c; Ding and Wang, 2020; Hu et al., 2021b), pose estimation (Jiang et al., 2020; Wang et al., 2019e, f; Ren et al., 2021, 2022; Zhou et al., 2023; Wang et al., 2024c, 2022b; Qian et al., 2023; Yan et al., 2024b), crowd counting (Zou et al., 2017, 2018b; Choi et al., 2022; Jiang et al., 2023; Xi et al., 2014), and more. Numerous survey papers have comprehensively summarized this flourishing research area, examining existing studies from various perspectives such as methodologies, application domains, and technical challenges (Tan et al., 2022; Hernandez and Bulut, 2022; Zhang et al., 2020a; Ma et al., 2019; Liu et al., 2019; Xiao et al., 2016; Li et al., 2021a; Jiang et al., 2018a; Yang et al., 2013; Yousefi et al., 2017).
Wi-Fi sensing systems face significant deployment challenges in real-world environments. This is primarily because Wi-Fi signals are intricately linked to their surroundings, so any alteration in the environment or the people within it can modify the signal characteristics. Even when the same person performs an identical action in the same setting, variations in position or orientation can result in inconsistent signal features, complicating reliable action recognition. To enable more widespread application of Wi-Fi sensing, researchers increasingly use advanced techniques like domain adaptation, domain generalization, and meta-learning to overcome these issues.
Recent studies have increasingly focused on enhancing the generalizability of Wi-Fi sensing techniques. For instance, Chen et al (Chen et al., 2023) adopts a cross-domain perspective, reviewing and categorizing over 60 papers that address challenges in Wi-Fi sensing. This work organizes various techniques—including domain-invariant feature extraction, transfer learning, and few-shot learning—that contribute to cross-domain robustness. In contrast, Wang et al (Wang et al., 2024a) zeroes in on the few-shot learning paradigm, providing a detailed overview of approaches such as transfer learning, metric learning, and meta-learning. Together, these studies highlight the significant progress made in applying advanced learning strategies to improve the adaptability and performance of Wi-Fi sensing systems.
Compared with previous works, this survey encompasses a significantly broader range of Wi-Fi generalization studies—over 200 papers in total. We have organized these studies along the entire Wi-Fi sensing pipeline: from device deployment and signal processing to feature learning and model deployment. This structured approach enables researchers to quickly locate the techniques that best address their specific challenges. Furthermore, we provide a detailed overview of open-source datasets, helping researchers efficiently identify the data needed to validate their algorithms. Finally, we offer a comprehensive analysis of future directions, including trends in multimodal approaches and large language models, to inspire new research ideas.
In summary, this work serves as a valuable resource for researchers seeking convenient access to methodologies, datasets, and innovative research strategies.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6em,align=left,fit=band,font=,, where level=2text width=8.2em,align=left,fit=band,font=,, where level=3text width=6.8em,font=,, where level=4text width=5em, font=,, [ Wi-Fi Human Sensing Generalization Categories, ver [ Hardware
Setup
2.1) , fill=orange!5 [ Distributed Antennas
2.1.1), fill=orange!10 [ WiNDR (Qin et al., 2024), WiCross (Qin et al., 2023), WiSDAR (Wang et al., 2018) , leaf,fill=orange!10, text width=40em ] ] [ Distributed Devices
2.1.2),fill=orange!10 [ ReWiS (Bahadori et al., 2022), Widar 3.0 (Zheng et al., 2019), OneFi (Xiao et al., 2021), WiTraj (Wu et al., 2021b) , leaf,fill=orange!10, text width=40em ] ] [ Scaling Up Dataset
2.1.3),fill=orange!10 [ XRF55 (Wang et al., 2024b), Widar 3.0 (Zheng et al., 2019), Zhang et al.(Zhang et al., 2023h), MM-Fi(Yang et al., 2024a) , leaf,fill=orange!10, text width=40em ] ] ] [ Signal
Preprocessing
2.2), fill=violet!5 [ Signal Analysis
Methods (§2.2.1), fill=violet!10 [ WiHand (Lu et al., 2019), CLAR (Zhou et al., 2022a), Gait-Enhance(Yang et al., 2023a), Zhuo et al.(Zhuo et al., 2023), Abuhoureyah et al. (Abuhoureyah et al., 2024), leaf, fill=violet!15, text width=40em ] ] [ Advanced Handcrafted
Indicators (§2.2.2), fill=violet!10 [ WiHGR (Meng et al., 2021), PhaseAnti (Huang et al., 2021), DPSense (Gao et al., 2022a), HandGest (Zhang et al., 2022d), Yu et.al. (Yu et al., 2023), AF-ACT (Zhang et al., 2022e), RoSeFi (Peng et al., 2023),
Wang et.al. (Wang et al., 2024e), WiGesFree(Ding et al., 2024), LiKey (Peng et al., 2024), Bu et.al. (Bu et al., 2018), Zhang et.al. (Zhang et al., 2023e) , leaf, fill=violet!15, text width=40em ] ] [ Motion Indicators
2.2.3), fill=violet!10 [ Sharp (Meneghello et al., 2022), OneSense (Zhao et al., 2024d), Yin et al.(Yin et al., 2021), Wi-learner (Feng et al., 2022), Niu et al. (Niu et al., 2021), and LAGER (Chen et al., 2024a), WiLife (Li et al., [n. d.]),
Mini-Batch Alignment (van Berlo et al., 2023, 2022), WiTraj (Wu et al., 2021b), Widar 3.0 (Zheng et al., 2019). Zhang et al. (Zhang et al., 2023i), WiTransformer (Yang et al., 2023c),
Shi et al. (Shi et al., 2023), Bulugu (Bulugu, 2023). GaitID (Zhang et al., 2020d), OneFi (Xiao et al., 2021), XFall (Chi et al., 2024), WiGesture (Gao et al., 2021), Shi et al. (Shi et al., 2020a),
Wang et al. (Wang et al., 2022c), Widir (Wu et al., 2016), Wang et al. (Wang et al., 2016e), WiNDR(Qin et al., 2024), WiDIGR (Zhang et al., 2019a), Wi-PIGR (Zhang et al., 2021d),
WiDetect (Zhang et al., 2019b), Zhu et al. (Zhu et al., 2024), leaf, fill=violet!15, text width=40em ] ] [ Angle Indicators
2.2.4), fill=violet!10 [ Li et al. (Li et al., 2020), EasyTrack (Han et al., 2023), 3D-ID (Ren and Yang, 2022), leaf, fill=violet!15, text width=40em ] ] ] [ Feature
Learning
2.3),fill=yellow!5 [ Domain Alignment
2.3.1),fill=yellow!10 [ JADA (Zou et al., 2018a), CrossGR (Li et al., 2021b), WiHARAN (Wang et al., 2021a), i-Sample (Zhou et al., 2024), CsiGAN (Xiao et al., 2019), CycleGAN (Zhu et al., 2017a), Wi-Adaptor (Zhang et al., 2021e),
GESFI (Zhang et al., 2024a), Khattak (Khattak and Khan, 2022), DINN (Zhou et al., 2021), INDG-Fi (Yang et al., 2024b), freeGait (Yan et al., 2024c), CATFSID (Wei et al., 2024), ADA (Zinys et al., 2021), DAFI (Li et al., 2021c),
Shi et al. (Shi et al., 2020a), DATTA (Strohmayer et al., 2024), WiSR (Liu et al., 2023a), WiCAR (Wang et al., 2019d), Berlo et al. (Berlo et al., 2023), WiCross (Qin et al., 2023), alignment.Li et al. (Li et al., 2021e),
Hassan et al. (Hassan et al., 2024), WiSDA (Jiao et al., 2024), LAGER (Chen et al., 2024a), WiCAU (Cui et al., 2024), CrossFi (Zhao et al., 2024a), P3ID (He et al., 2024), AdapLoc (Zhou et al., 2020), WiAi-ID (Liang et al., 2023),
Chen et al. (Mehryar, 2023), Zhang et al. (Zhang et al., 2021b), Mini-Batch Alignment (van Berlo et al., 2023, 2022), MetaGanFi (Zhang et al., 2022b), CsiGAN (Xiao et al., 2019), cyclegan (Zhu et al., 2017a),
SIDA (Zhang et al., 2023b), Zhang et al. (Zhang et al., 2022a), EI (Jiang et al., 2018c), AirFi (Wang et al., 2022d)CrossID (Wu et al., 2021a), Wang et al. (Wang et al., 2024d), Wi-Piga (Hao et al., 2021), Wi-Dist (Zhang et al., 2024d),
UH-Sense (Chen et al., 2024b), Person-in-WiFi (Wang et al., 2019f), Elujide et al. (Elujide et al., 2022), SCTRFL (Zheng et al., 2020), PIAS (Xiao et al., 2024b), WiLISensing (Ding et al., 2020b) , leaf,fill=yellow!15, text width=40em ] ] [ Meta-learning
2.3.3),fill=yellow!10 [ MAML (Finn et al., 2017), Metaloc (Gao et al., 2023), RoMF (Zhang et al., 2024b), WiLiMetaSensing (Ding et al., 2021a), MetaGANFi (Zhang et al., 2022b), WiOpen (Zhang et al., 2024c), CSI-GDAM (Zhang et al., 2021a),
FewCS (Yang et al., 2023b), WiGesID (Zhang et al., 2022c), MatNet-eCSI (Shi et al., 2020b), LT-WiOB (Zhou et al., 2022c), TransferSense (Bu et al., 2022), DAFI  (Li et al., 2021c), MetaGANFi (Zhang et al., 2022b),
CLAR-Xiao et al. (Xiao et al., 2024a), DualConFi (Xu et al., 2022), SFTSeg  (Xiao et al., 2022), HAR-MN-EF (Shi et al., 2020c), DA-HAR (Sheng et al., 2023), Shi et al. (Wang et al., 2022e), WiDeo(Islam and Nirjon, 2021)
RF-Net (Ding et al., 2020a), Wi-learner (Feng et al., 2022), Reptile (Nichol, 2018), ML-WiGR (Gao et al., 2022b), TOSS (Zhou et al., 2022b), Zhang et al. (Zhang et al., 2020b), Huang et al. (Huang et al., 2022),
AFSL-HAR (Wang et al., 2021b), Wi-AM (Zhang et al., 2023j), Prototypical Network (Snell et al., 2017), MetaFormer (Sheng et al., 2024) , leaf,fill=yellow!15, text width=40em ] ] [ Embedding Learning
2.3.4),fill=yellow!10 [ FewCS (Yang et al., 2023b), ReWiS (Bahadori et al., 2022), WiGesID  (Zhang et al., 2022c), MatNet-eCSI (Shi et al., 2020b), LT-WiOB (Zhou et al., 2022c), TransferSense  (Bu et al., 2022), 3D-ID  (Ren and Yang, 2022),
WiDeo (Islam and Nirjon, 2021), CrossFi  (Zhao et al., 2024a), SFTSeg  (Xiao et al., 2022), CSI-GDAM (Zhang et al., 2021a), UniFi (Liu et al., 2024b), MaP-SGAN (Liang et al., 2024), Wi-Cro (Mao et al., 2024), CATFSID (Wei et al., 2024),
WiOpen (Zhang et al., 2024c), WiGR (Hu et al., 2021a), DAFI (Li et al., 2021c), Chen et al. (Chen and Chang, 2022), Ding et al. (Ding et al., 2021b), WiONE (Gu et al., 2021), Shi et al. (Shi et al., 2022), Caution (Wang et al., 2022e),
LT-WiOB (Zhou et al., 2022d), Wang et al. (Wang et al., 2020), Zhou et al. (Zhao et al., 2024c), CSI-GLSTN (Zhang et al., 2024e), HAR-MN-EFF (Shi et al., 2020c), DA-HAR (Sheng et al., 2023), CFLH (Wang et al., 2024f),
DualConFi (Xu et al., 2022), CLAR-Xiao et al. (Xiao et al., 2024a), CrossSense (Zhang et al., 2018), Sugimoto et al. (Sugimoto et al., 2023), CSI-MTGN (Zhang et al., 2023a), Wi-Fringe (Islam and Nirjon, 2020),
WiFi2Radar (Nirmal et al., 2024), i-Sample (Zhou et al., 2024) , leaf,fill=yellow!15, text width=40em ] ] [ Data augmentation
2.3.5),fill=yellow!10 [ Wi-Dist (Zhang et al., 2024d), WiARGAN (Huang et al., [n. d.]), WiSGP (Liu et al., 2023b), Wang et al. (Wang et al., 2020), WGAN (Arjovsky et al., 2017), LESS (Zhang et al., 2023g), WiAG (Virmani and Shahzad, 2017), AirFi (Wang et al., 2022d),
CrossGR (Li et al., 2021b), Fido (Chen et al., 2020), Fidora (Chen et al., 2022), RFBoost (Hou and Wu, 2024), CLAR-Xiao et al. (Xiao et al., 2024a), Wi-Cro (Mao et al., 2024), freeGait (Yan et al., 2024c), SFTSeg (Xiao et al., 2022),
WiTeacher (Xiao et al., 2023), adversarial autoencoder, (AAE) (Makhzani et al., 2015), freeLoc (Yan et al., 2024a) , leaf,fill=yellow!15, text width=40em ] ] [ Pseudo-labeling and
pure feature learning
2.3.6),fill=yellow!10 [ LAGER (Chen et al., 2024a), P3ID (He et al., 2024), OpenFi (Liu et al., 2024a), CLAR (Zhou et al., 2022a), TOSS (Zhou et al., 2022b), MAML (Finn et al., 2017), Zhang et.al (Zhang et al., 2023j), DA-HAR (Sheng et al., 2023), PIAS (Xiao et al., 2024b),
PAC-CSI (Su et al., 2023), 3D-ID (Ren and Yang, 2022), Abuhoureyah et al. (Abuhoureyah et al., 2023), Ma et al. (Ma et al., 2021), (Sheng et al., 2020) , leaf,fill=yellow!15, text width=40em ] ] ] [ Deployment
2.4), fill=teal!3 [ Transfer Learning
2.4.1), fill=teal!7 [ XRF55 (Wang et al., 2024b), BullyDedect (Lan et al., 2024), WiLISensing (Ding et al., 2020b), DADA-AD (Chen et al., 2021), Brinke & Meratnia (Brinke and Meratnia, 2019b), WiTransfer (Fang et al., 2020),
FewSense (Yin et al., 2022), ResMon (Zheng et al., 2023), RoGER (Bi et al., 2024), Hou et al. (Hou et al., 2022a), DASECount (Hou et al., 2022b) , leaf, fill=teal!10,, text width=40em ] ] [ Federated Learning
2.4.2), fill=teal!7 [ WiFederated (Hernandez and Bulut, 2021), pFedBKD (Geng et al., 2023), CARING (Li et al., 2023), Qi et al. (Qi et al., 2023), Co-WiSensing (Zhang et al., 2023d), Zhang et al. (Zhang et al., 2021b),
AdaWiFi (Zheng et al., 2024), Zhang et al. (Zhang et al., 2023f) , leaf, fill=teal!10,, text width=40em ] ] [ Continuous Sensing
2.4.3), fill=teal!7 [ FewCS (Yang et al., 2023b), WiDrive (Bai et al., 2019), M-WiFi (Soltanaghaei et al., 2020), RISE (Zhai et al., 2021), CCS (Fu et al., 2024), CSI-ARIL (Zhang et al., 2023c) , leaf, fill=teal!10, text width=40em ] ] ] ]
Figure 1. Categories of Cross-Domain Wi-Fi Human Sensing.
2. Taxonomy
Achieving generalizability in Wi-Fi sensing involves addressing challenges across different stages, including device deployment, signal processing, feature learning, and model deployment. While some studies focus on a single stage, others span multiple stages, tackling two or even three phases simultaneously. To provide a clear and systematic overview of existing research, we organize the surveyed works along the Wi-Fi sensing pipeline, covering every step from device setup to model deployment.
2.1. Experimental Setup
At the experimental setup stage, some studies enhance the generalization capability of Wi-Fi sensing by adopting distributed antenna deployment, integrating additional devices, and collecting more extensive datasets.
2.1.1. Distributed Antennas
WiNDR (Qin et al., 2024) and WiCross (Qin et al., 2023) position three antennas from the transmitter and three antennas from the receiver in a staggered, distributed manner around a 360-degree circle, with the subject performing gestures at the center. This configuration achieves direction-agnostic gesture recognition. Additionally, WiSDAR (Wang et al., 2018) tests various distributed antenna topologies, including line, hexagon, square, and random shapes, to explore regions that are most sensitive to human activities, achieving highly accurate and reliable recognition results.
2.1.2. Distributed Devices
Some studies utilize additional WiFi devices to enhance sensing range and improve generalization to human orientation and position. For instance, ReWiS (Bahadori et al., 2022), Widar 3.0 (Zheng et al., 2019), and OneFi (Xiao et al., 2021) adopt configurations of 1 transmitter with 6 receivers and 1 transmitter with 4 receivers, respectively, to extract environment-independent features. WiTraj (Wu et al., 2021b) leverages 3 receivers placed at different viewing angles to capture human walking and achieve robust trajectory reconstruction. Additionally, Wang et al. (Wang et al., 2022a) propose a theoretical model analyzing how the distance between the transmitter and receiver influences the sensing range. Optimizing the placement of these devices can effectively enhance the sensing area and reduce the impact of environmental interference.
2.1.3. Scaling up dataset
Scaling up training datasets is another effective approach to enhancing generalization capability, as demonstrated in Widar 3.0 (Zheng et al., 2019), Zhang et al. (Zhang et al., 2023h), MM-Fi(Yang et al., 2024a), and XRF55 (Wang et al., 2024b). For example, XRF55 (Wang et al., 2024b) shows that training on approximately 30,000 samples enables the sensing model to naturally achieve direction and position invariance, while also supporting adaptation to new environments through few-shot fine-tuning.
2.2. Signal Preprocessing Stage
Many works focus on the signal preprocessing stage to extract domain-invariant features to enhance the generalizability of Wi-Fi sensing. These approaches can be broadly categorized into four types: (1) using signal analysis methods, such as Principal Component Analysis (PCA), to extract action-related features from raw signals; (2) processing phase or amplitude information to generate advanced action indicators; (3) extracting motion-specific indicators like Doppler Frequency Shift (DFS) and Body Velocity Profile (BVP) to capture dynamic movements; (4) employing antenna array techniques to derive measurements such as Angle of Arrival (AoA) and Angular Difference of Arrival (ADoA) for sensing. Next, we will introduce the papers corresponding to these four categories.
2.2.1. Signal Analysis Methods
Signal analysis plays a crucial role in Wi-Fi sensing by extracting relevant features from raw signals, ensuring generalizability across various environments, locations, and orientations. WiHand (Lu et al., 2019) employs a Low Rank and Sparse Decomposition (LRSD) algorithm to separate gesture signals from background noise, enhancing its resilience to location variations. CLAR (Zhou et al., 2022a) utilizes Singular Spectrum Analysis (SSA) (Golyandina et al., 2001) to extract the trend components from Wi-Fi signals, enabling robust activity recognition. Gait-Enhance (Yang et al., 2023a) introduces a CSI reconstruction algorithm that mitigates the negative impact on recognition accuracy caused by varying CSI patterns associated with walking in different directions. Zhuo et al.  (Zhuo et al., 2023) apply a combination of Principal Component Analysis (PCA) and Variational Mode Decomposition (VMD) (Dragomiretskiy and Zosso, 2013) to reduce noise interference in non-ideal sleep positions, enabling position-free breath detection. Abuhoureyah et al. (Abuhoureyah et al., 2024) leverage Independent Component Analysis (ICA) and Continuous Wavelet Transform (CWT) to decompose CSI signals, improving multi-user human action recognition in complex environments.
2.2.2. Advanced Handcrafted Indicators
Many researchers have focused on exploring robust handcrafted features for domain-independent sensing. WiHGR (Meng et al., 2021) uses a Phase Difference Matrix to extract phase-related features. PhaseAnti (Huang et al., 2021) introduces Nonlinear Phase Error Variation (NLPEV), which is independent of Cochannel Interference (CCI), to mitigate interference and enhance recognition accuracy. DPSense (Gao et al., 2022a) proposes the Error of Dynamic Phase Index (EDP-index) to evaluate the sensing quality of signal segments, prioritizing high-quality segments to improve gesture recognition performance. HandGest (Zhang et al., 2022d) combines Dynamic Phase Vector (DPV) and Motion Rotation Variable (MRV) for in-the-air handwriting recognition. Similarly, Yu et al. (Yu et al., 2023) extract dynamic phase features for position-independent gesture recognition. Additionally, AF-ACT (Zhang et al., 2022e) extracts semantic activity and temporal features to characterize activities across various locations. RoSeFi (Peng et al., 2023) identifies contextual association and waveform mirror symmetry in CSI data, introducing the Waveform Mirror Symmetry Factor (WMSF) to quantify symmetry between sit-stand postural transition (SPT) CSI data, aiding in sedentary behavior detection.
Some works also extract features in the IQ complex plane. For instance, Wang et al. (Wang et al., 2024e) analyze the curvature of signal curves over time in the IQ plane. WiGesFree(Ding et al., 2024) introduces the Sample Point Distance (SPD), which uses changing CSI ratio patterns in the complex plane to capture gesture information accurately. LiKey (Peng et al., 2024) leverages rotational direction features in the CSI complex plane as an extraction of location-independent features. Other works transform Wi-Fi CSI data into images and use visual methods, such as pre-trained models and data augmentation, to improve generalization capabilities. Bu et al. (Bu et al., 2018) convert the amplitude of each CSI stream into a grayscale CSI image and fine-tune pre-trained networks like VGG16 and VGG19 for feature extraction. Zhang et al. (Zhang et al., 2023e) convert one-dimensional CSI time series into two-dimensional recurrence plots (RP) and apply data augmentation techniques like horizontal flipping, cropping, and color distortion to improve model performance and robustness in action recognition.
2.2.3. Motion Indicators
Many works compute domain-independent motion indicators in speed, velocity, direction, and trajectory, to enhance the generalizability of Wi-Fi sensing systems. Sharp (Meneghello et al., 2022), OneSense (Zhao et al., 2024d), Yin et al. (Yin et al., 2021), Wi-learner (Feng et al., 2022), Niu et al. (Niu et al., 2021), and LAGER (Chen et al., 2024a) focus on utilizing Doppler Frequency Shift (DFS). Mini-Batch Alignment (van Berlo et al., 2023, 2022) incorporates both DFS and Gramian Angular Difference Field (GADF). WiLife (Li et al., [n. d.]) further explores Doppler velocity as a key motion indicator. WiTraj (Wu et al., 2021b) identifies speed ambiguity in the Doppler-MUSIC algorithm’s spectrum. To address this, it employs multiple receivers at different viewing angles and leverages the CSI quotient to eliminate this ambiguity, enabling robust trajectory reconstruction.
Widar 3.0 (Zheng et al., 2019) introduces the Body-coordinate Velocity Profile (BVP), which captures the power distribution across different velocities of body parts involved in gesture movements. This approach enhances generalization to new locations, environments, and subjects while maintaining performance consistency. BVP and its variants have demonstrated remarkable generalization capabilities. For instance, Zhang et al. (Zhang et al., 2023i) and WiTransformer (Yang et al., 2023c) adopt BVP, while Shi et al. (Shi et al., 2023) and Bulugu (Bulugu, 2023) extend it to 3D BVP. GaitID (Zhang et al., 2020d) utilizes gait-specific BVP (gait-BVP), OneFi (Xiao et al., 2021) employs velocity distribution, and XFall (Chi et al., 2024) incorporates a speed distribution profile to enhance robustness. In addition to BVP, WiGesture (Gao et al., 2021) proposes the Motion Navigation Primitive (MNP) to improve gesture recognition accuracy. Shi et al. (Shi et al., 2020a) convert CSI amplitude and relative amplitude into frequency-domain representations, such as STFT holograms, to characterize the moving speeds of different body parts. Wang et al. (Wang et al., 2022c) use phase changes of the reflection path to relate the speed vector and gesture trajectory in the body coordinate system, implicitly capturing the impact of position and orientation, which is estimated by a preamble gesture in each new environment.
The Fresnel Zone Model in physics is also widely used to enhance the generalization capability of Wi-Fi sensing. Widir (Wu et al., 2016) utilizes it to capture walking direction. Wang et al.(Wang et al., 2016e) apply the model to detect breast motion for respiration detection. WiNDR(Qin et al., 2024) leverages the model for direction-agnostic gesture recognition. Additionally, WiDIGR (Zhang et al., 2019a) and Wi-PIGR (Zhang et al., 2021d) use the Fresnel Zone Model to eliminate direction information for gait recognition. Statistical EM methods have also been applied to Wi-Fi sensing. WiDetect (Zhang et al., 2019b) uses statistical EM theory to model motion detection, connecting the Autocorrelation Function (ACF) of CSI to target motion, independent of environment, location, orientation, and subjects. Zhu et al. (Zhu et al., 2024) extract the Amplified ACF spectrum from CSI, enabling motion features to be isolated, independent of environmental factors.
2.2.4. Angle Indicators
Angle indicators are also used to enhance the generalizability of Wi-Fi sensing. Li et al. (Li et al., 2020) use Angle Difference of Arrival (ADoA) for location-free CSI-based activity recognition. EasyTrack (Han et al., 2023) utilizes AoA–AoD values, leveraging changes in the length and direction of the Wi-Fi signal path. Through incremental calculations and dynamic path extraction, it can reconstruct gesture trajectories in real-time without relying on the initial gesture position or a fixed device location. 3D-ID (Ren and Yang, 2022) uses 2D AoA estimation of signal reflections to enable Wi-Fi to visualize a person in the environment. This visualization is then digitized into a 3D body representation using deep learning, which extracts both the static body shape and dynamic walking patterns for person re-identification.
2.3. Feature Learning Stage
The Feature Learning Stage is fundamental to improving the adaptability and robustness of Wi-Fi sensing systems across diverse domains. This stage includes several key techniques aimed at overcoming the challenges of domain variability, limited labeled data, and environmental changes. Methods such as domain alignment, reconstruction structures, embedding learning, data augmentation, pseudo-labeling and pure feature learning work together to enhance model performance, ensuring that features learned from one domain can be successfully applied to others. These approaches through techniques like adversarial training, similarity metrics, data synthesis, and embedding refinement, enable Wi-Fi sensing models to generalize effectively in dynamic and heterogeneous real-world environments.
2.3.1. Domain Alignment
Domain alignment techniques are crucial for reducing the gap between source and target domains in Wi-Fi sensing systems. These methods ensure better generalization by aligning feature distributions, enabling models to adapt to new environments. In the domain alignment with domain discriminator approach, JADA (Zou et al., 2018a) applies domain adversarial training to classify features as belonging to either the source or target domain, facilitating robust feature alignment. CrossGR (Li et al., 2021b) uses a GAN to generate synthetic data and employs a discriminator to distinguish between real and synthetic samples, ensuring that domain alignment is achieved effectively. Similarly, WiHARAN (Wang et al., 2021a) incorporates a domain discriminator to align activity features across different environments, thus enhancing activity recognition. In a similar vein, i-Sample (Zhou et al., 2024) augments datasets through gradient learning while utilizing a discriminator to identify and align both real and synthetic samples, improving the adaptability of the model. Zhang et al. (Zhang et al., 2021b) employs a feature encoding network, which is deployed in the local server of each environment. The Johnson-Lindenstrauss transform is applied to preserve privacy information before it is sent to the central server. An environment classifier distinguishes which environment the input datum comes from, and as the system converges, it becomes environment-invariant, making it incapable of distinguishing between environments.
2.3.2. Adversarial Domain Adaptation
Adversarial Domain Adaptation (Ganin et al., 2016; Tzeng et al., 2017) aims to align the source and target domains with adversarial training, enabling Wi-Fi sensing to generalize across new individuals, environments, and other novel domains. For example, JADA (Zou et al., 2018a) feeds source-domain data into a source-domain encoder and target-domain data into a target-domain encoder. Through adversarial training with a domain discriminator, the features extracted by both encoders become indistinguishable regarding their domain of origin. The source-domain encoder is then frozen, and a classification head is trained using labeled source-domain data for action recognition. Finally, the target-domain encoder and classification head are combined for inference on target-domain data, achieving cross-domain action recognition. CrossGR (Li et al., 2021b) trains an eliminator to make action-related features unclassifiable, isolating components irrelevant to the actions. By removing these irrelevant components, Wi-Fi features are better suited for cross-person and cross-environment action recognition. In Person-in-WiFi (Wang et al., 2019f), pairs of data inputs are used to pre-train a discriminator that determines whether the data pairs come from the same environment. A generator is then trained to ensure that the features of these pairs are consistently judged as originating from the same environment. This process removes environmental information and enhances the accuracy of cross-scene human pose estimation.
Domain Alignment with Domain Classifier approaches explicitly classify features into domains and use this feedback for alignment. EI (Jiang et al., 2018c) employs a domain classifier with the objective of minimizing the loss function to maximize the performance of domain label prediction, which contrasts with the ultimate goal of learning domain-independent features for activities. CsiGAN (Xiao et al., 2019), supported by CycleGAN (Zhu et al., 2017a), generates virtual samples and employs a domain classifier to distinguish real and synthetic data for cross-domain adaptation. Wi-Adaptor (Zhang et al., 2021e) uses domain classifiers and adversarial training to align features across source and target domains. GESFI (Zhang et al., 2024a) and Khattak (Khattak and Khan, 2022) incorporate domain classifiers to address variability in gesture recognition and location-based sensing, respectively. DINN (Zhou et al., 2021) extracts domain-invariant features for pose estimation, while INDG-Fi (Yang et al., 2024b) and freeGait (Yan et al., 2024c) utilize domain randomization and adversarial techniques to ensure robust alignment. CATFSID (Wei et al., 2024), ADA (Zinys et al., 2021), and DAFI (Li et al., 2021c) combine domain classifiers with techniques like triplet losses and gesture discriminators to align features and improve cross-domain generalization. Furthermore, Shi et al. (Shi et al., 2020a) and DATTA (Strohmayer et al., 2024) focus on dynamic alignment using domain classifiers for test-time adaptation. Other works such as WiSR (Liu et al., 2023a), WiCAR (Wang et al., 2019d), and Berlo et al. (Berlo et al., 2023) explore multi-domain classifiers for cross-device and cross-environment sensing, while WiCross (Qin et al., 2023) uses classifiers for multi-environment adaptation.
Domain Alignment with Similarity Matrix employs similarity metrics such as Maximum Mean Discrepancy (MMD) or cosine similarity for feature alignment. Li et al. (Li et al., 2021e), Hassan et al. (Hassan et al., 2024) and WiSDA (Jiao et al., 2024) leverage cosine similarity to align source and target features, while LAGER (Chen et al., 2024a) and WiCAU (Cui et al., 2024) apply MMD and Wasserstein distance for distribution alignment. CrossFi (Zhao et al., 2024a) uses Siamese networks and MK-MMD to align distributions for few-shot and zero-shot adaptation. P3ID (He et al., 2024) combines DTW-based soft labeling with MMD, and AdapLoc (Zhou et al., 2020) aligns probability distributions using Euclidean distances. Other methods, such as WiAi-ID (Liang et al., 2023), Chen et al. (Mehryar, 2023) , and Mini-Batch Alignment (van Berlo et al., 2023, 2022), use advanced similarity metrics for fine-grained alignment. AirFi (Wang et al., 2022d) aligns multiple source domains using MMD and then applies it to the target domain by extracting common features from various training environments and minimizing the distribution differences between them. An arbitrary Gaussian noise is added to each CSI data sequence, generating new CSI datasets that help approximate the CSI data distribution in different environments, which benefits system training.
Domain Alignment with GAN focuses on generating synthetic samples for better alignment. MetaGanFi (Zhang et al., 2022b) integrates CycleGAN for domain alignment while leveraging meta-learning to enhance generalization.CsiGAN (Xiao et al., 2019) uses cyclegan (Zhu et al., 2017a), SIDA (Zhang et al., 2023b) uses VAE and GANto generate domain-specific data and align minority domains to the majority. Zhang et al. (Zhang et al., 2022a) uses CycleGAN to convert source features into target domain representations for improved cross-domain performance. WiLISensing (Ding et al., 2020b) first uses MMD to align multiple source domains and then leverages transfer learning to transfer features between different locations, freezing layers before the fully connected layers and training with a small number of samples to enhance model adaptability. PIAS (Xiao et al., 2024b) designs a cluster to generate pseudo-labels for target domain features and filters out unreliable training samples, using Contrastive Domain Discrepancy (CDD) loss to minimize intra-class feature discrepancies and maximize inter-class feature discrepancies. SCTRFL (Zheng et al., 2020) utilizes Restricted Boltzmann Machines (RBM) for amplitude and phase reconstruction in new environments. Distorted CSI amplitude and processed phase are input into trained autoencoders, which rebuild calibrated CSI from real-time measurements, restoring them to their reference state. Component separation with a discriminator is used in works like Elujide et al. (Elujide et al., 2022), which applies unsupervised adversarial invariance, and Person-in-WiFi (Wang et al., 2019f). UH-Sense (Chen et al., 2024b) uses GAN to remove noise from features, while Wi-Dist (Zhang et al., 2024d) uses CycleGAN to generate more data, training the classifier and feature generator adversarially to mitigate the negative impact of environmental changes. Wi-Piga (Hao et al., 2021) addresses personnel independence in action recognition using GANs, while Wang et al. (Wang et al., 2024d) decouples features related to identity and activity with separate task classifiers. In the domain-specific memory approach, CrossID (Wu et al., 2021a) learns affine parameters and identity attributes modeled as the interrelationship of affine parameters, which can be conditioned through instance normalization.
2.3.3. Meta-learning
Meta-learning techniques are designed to enable Wi-Fi sensing models to quickly adapt to new domains or tasks with minimal labeled data. These approaches are especially valuable when dealing with scenarios where training data is scarce or environments are highly dynamic. In the case of MetaFormer (Sheng et al., 2024), meta-learning is applied through the Prototypical Network (Snell et al., 2017), which learns class prototypes for few-shot classification tasks. MetaGanFi (Zhang et al., 2022b) uses CycleGAN to align multiple domains into a unified domain and integrates MAML (Finn et al., 2017) for efficient meta-learning, allowing the model to quickly adapt to new domains. WiLiMetaSensing (Ding et al., 2021a) adopts a matching network to learn domain-invariant representations by matching similar data points from different domains. Wi-AM (Zhang et al., 2023j) uses meta-learning techniques to optimize activity recognition, although the specific meta-learning method is not detailed. AFSL-HAR (Wang et al., 2021b) employs MAML (Finn et al., 2017) to adapt human activity recognition models across domains with minimal labeled data. RoMF (Zhang et al., 2024b) combines meta-learning with Graph Neural Networks (GNNs) for cross-domain activity recognition. The specific meta-learning algorithm is not identified, but the combination of graph-based learning with meta-learning enhances model performance across environments. Huang et al. (Huang et al., 2022) uses MAML (Finn et al., 2017) for efficient adaptation in localization tasks, enabling fast adaptation to new environments. Metaloc (Gao et al., 2023) uses MAML (Finn et al., 2017) to adapt localization models in new environments with minimal data, improving generalization to unseen domains. Zhang et al. (Zhang et al., 2020b) propose a meta-learning approach that enables models to be fine-tuned with very little training effort in new environments or for recognizing new types of activities, based on the Learning to Learn principle (Andrychowicz et al., 2016). In TOSS (Zhou et al., 2022b), a dynamic pseudo-labeling strategy is combined with MAML for few-shot learning. The model uses both labeled and unlabeled target samples, leveraging MAML to adapt quickly to new domains. ML-WiGR (Gao et al., 2022b) uses Reptile (Nichol, 2018), another meta-learning algorithm, to improve the generalization of Wi-Fi sensing systems across different domains. Similarly, Wi-learner (Feng et al., 2022) employs MAML and Doppler Frequency Shift (DFS) for rapid domain adaptation in human activity recognition tasks. Finally, RF-Net (Ding et al., 2020a) uses metric-based meta-learning for one-shot human activity recognition, enabling the model to recognize new activities with only a single labeled example from each class.
2.3.4. Embedding Learning
Embedding learning focus on constructing compact, robust, and domain-invariant feature representations to enhance the adaptability of Wi-Fi sensing systems in diverse domains. These methods are especially useful for scenarios like few-shot learning, zero-shot learning, and cross-domain recognition. Some approaches leverage few-shot and prototypical learning frameworks, such as FewCS (Yang et al., 2023b) and ReWiS (Bahadori et al., 2022), which utilize prototypical networks to learn class prototypes for robust cross-domain recognition. Similarly, WiGesID  (Zhang et al., 2022c) employs relation networks to model the relationships between input samples and improve the generalization of domain-invariant embeddings, while MatNet-eCSI (Shi et al., 2020b) integrates matching networks for effective few-shot learning in varying domains. LT-WiOB (Zhou et al., 2022c) and TransferSense  (Bu et al., 2022) use triplet losses to generate embeddings that preserve inter-class separation and intra-class compactness, enabling better performance in unseen environments. For re-identification and activity recognition, 3D-ID  (Ren and Yang, 2022) uses 2D AoA estimations and Siamese networks to construct embeddings for 3D body representations, enabling robust person re-identification across domains. Similarly, WiDeo (Islam and Nirjon, 2021) combines Wi-Fi and visual data to generate joint embeddings that enhance activity recognition across domains, while CrossFi  (Zhao et al., 2024a) integrates Siamese networks with MK-MMD for zero-shot and few-shot domain adaptation. Other embedding learning approaches rely on graph-based learning and pretext tasks. SFTSeg  (Xiao et al., 2022) applies time-series-specific pretext tasks and triplet loss to train models with limited labeled samples, while CSI-GDAM (Zhang et al., 2021a) leverages Graph Neural Networks (GNNs) and relation networks to create domain-invariant activity representations.
Some methods focus on integrating multi-view or triplet embeddings. For example, UniFi (Liu et al., 2024b) employs spatio-temporal multi-view networks with attention mechanisms and triplet loss to align multi-device embeddings, while MaP-SGAN (Liang et al., 2024) uses GANs to augment data and train Siamese networks for robust cross-domain embeddings. Similarly, Wi-Cro (Mao et al., 2024) combines triplet loss with synthetic data generated by CycleGAN for cross-domain adaptation. Loss function-based embedding approaches, such as CATFSID (Wei et al., 2024), WiOpen (Zhang et al., 2024c), and WiGR (Hu et al., 2021a), also utilize triplet losses to align source and target domain embeddings, ensuring robustness against domain-specific noise. In the context of matching networks and meta-learning, DAFI (Li et al., 2021c) uses matching networks with triplet losses to enhance cross-domain feature alignment. Other methods, like Chen et al. (Chen and Chang, 2022), Ding et al. (Ding et al., 2021b), and WiONE (Gu et al., 2021), apply matching and prototypical networks for cross-domain authentication and activity recognition, while Shi et al. (Shi et al., 2022), Caution (Wang et al., 2022e), and LT-WiOB (Zhou et al., 2022d) improve few-shot generalization through similar frameworks. GAN-based methods like Wang et al. (Wang et al., 2020) and Zhou et al. (Zhao et al., 2024c) integrate GAN-generated samples into prototypical embeddings to improve cross-domain recognition performance. Furthermore, methods such as CSI-GLSTN (Zhang et al., 2024e) use GNNs to learn hidden activity graph representations, enabling robust cross-domain activity detection, while HAR-MN-EFF (Shi et al., 2020c) and DA-HAR (Sheng et al., 2023) enhance embeddings for human activity recognition through matching networks and triplet loss.
Contrastive and multi-task learning are also widely adopted in embedding learning. CFLH (Wang et al., 2024f) applies contrastive learning to align embeddings by minimizing intra-class variance and maximizing inter-class separability. DualConFi (Xu et al., 2022) extends contrastive learning by generating additional samples to improve embedding robustness, while CLAR-Xiao et al. (Xiao et al., 2024a) uses a diffusion model for data augmentation combined with contrastive learning for embedding enhancement. Multi-task learning approaches, such as CrossSense (Zhang et al., 2018) and Sugimoto et al. (Sugimoto et al., 2023), train task-specific embeddings with shared parameters to create environment-invariant representations, while CSI-MTGN (Zhang et al., 2023a) aligns task-specific embeddings with domain-invariant features for improved generalization. Finally, cross-modal embedding techniques, like Wi-Fringe (Islam and Nirjon, 2020), integrate BERT embeddings for open-set action recognition by combining text and Wi-Fi signal modalities, while WiFi2Radar (Nirmal et al., 2024) translates Wi-Fi spectrograms into mmWave radar representations for robust cross-modal learning. Methods such as i-Sample (Zhou et al., 2024) combine embedding techniques with dataset augmentation, ensuring generalization across varying domains. Overall, embedding learning methods enhance the adaptability and robustness of Wi-Fi sensing systems by creating compact and domain-agnostic feature representations suitable for diverse applications.
2.3.5. Data augmentation
Data augmentation techniques play a crucial role in enhancing the generalization ability of Wi-Fi sensing models by generating synthetic data. These methods are particularly valuable when labeled data is scarce or environments are highly dynamic. freeLoc (Yan et al., 2024a) employs an adversarial autoencoder (AAE) (Makhzani et al., 2015) to augment source domain samples, using a domain classifier to recognize different users and motions, thus improving domain adaptability. Similarly, WiTeacher (Xiao et al., 2023) uses a StyleGAN to generate synthetic data from unlabeled target samples, adjusting labels dynamically with an adaptive label smoothing technique inspired by soft target methods (Müller et al., 2019; Xie et al., 2021). SFTSeg (Xiao et al., 2022) enhances limited labeled target samples by using triplet loss, proposing a time-series-specific pretext task that builds positive and negative sample pairs from unlabeled data, improving network training. FreeGait (Yan et al., 2024c) uses AAE for generating synthetic data that resembles labeled data, incorporating user IDs and jointly training with a domain classifier to align source and target domains. Wi-Cro (Mao et al., 2024) utilizes cyclyeGAN (Zhu et al., 2017a) to generate synthetic target domain data, applying triplet loss to improve generalization across domains. CLAR-Xiao et al. (Xiao et al., 2024a) propose a diffusion model for augmenting data, combining contrastive learning to learn embeddings. This model augments time-series data by dividing high- and low-frequency components to avoid distortion, ensuring better data quality. RFBoost (Hou and Wu, 2024) employs physical data augmentation techniques, including frequency-domain and time-domain data augmentation, as well as test-time augmentation, to improve model robustness across varying environmental conditions. Fidora (Chen et al., 2022) and Fido (Chen et al., 2020) use Variational Autoencoders (VAE) for augmenting labeled data, incorporating reconstruction branches to train a domain-adaptive classifier that adjusts to newly collected unlabeled data. CrossGR (Li et al., 2021b) leverages GAN to generate fine-grained synthetic data, with a domain discriminator distinguishing real and fake data, enhancing cross-domain generalization. AirFi (Wang et al., 2022d) uses MMD to align multiple source domains and applies feature augmentation to improve robustness across environments by adding Gaussian noise to the CSI data sequence. WiAG (Virmani and Shahzad, 2017) requests training samples for gestures in a single configuration and generates virtual samples for all possible configurations through a novel transformation function, enabling more robust learning. LESS (Zhang et al., 2023g) employs WGAN (Arjovsky et al., 2017) with a gradient penalty (Antoniou et al., 2018) and Relation Network (Sung et al., 2018) to generate diverse synthetic data. Wang et al. (Wang et al., 2020) integrate GAN (Goodfellow et al., 2014) for generating virtual training samples and use a Relation network (Sung et al., 2018) to enhance feature representation. WiSGP (Liu et al., 2023b) uses domain gradient to augment data, improving model generalization across different environments. WiARGAN (Huang et al., [n. d.]) utilizes GAN to expand and generate data, while Wi-Dist (Zhang et al., 2024d) employs CycleGAN to generate more data and trains both a classifier and feature generator adversarially, mitigating environmental changes and identifying suspicious activities.
2.3.6. Pseudo-labeling and pure feature learning
Pseudo-labeling techniques have gained significant attention for leveraging unlabeled target domain data to improve model performance. Methods like LAGER (Chen et al., 2024a) use a pre-heated strategy for pseudo-labeling target domain data, aligning the estimated mean features between source and target domains. P3ID (He et al., 2024) applies DTW similarity and iterative soft labeling to annotate unlabeled data, using MMD for alignment between domains. OpenFi (Liu et al., 2024a) and CLAR (Zhou et al., 2022a) further enhance this process by generating pseudo-labels using virtual embeddings and multi-classifiers, respectively. CLAR (Zhou et al., 2022a) employs Singular Spectrum Analysis (SSA) (Golyandina et al., 2001) to extract trend components from activity samples and then annotates them with pseudo-labels for activity recognition. TOSS (Zhou et al., 2022b) introduces a dynamic pseudo-label strategy using unlabeled target samples and labeled samples, coupled with MAML (Finn et al., 2017) for fast adaptation. Zhang et.al (Zhang et al., 2023j) propose pseudo-labeling combined with consistency regularization to eliminate domain feature discrepancies, while DA-HAR (Sheng et al., 2023) applies triplet loss and pseudo-labeling for target domain adaptation. PIAS (Xiao et al., 2024b) generates pseudo-labels through clustering and filters unreliable samples to minimize domain discrepancies using Contrastive Domain Discrepancy (CDD) loss. These methods enable the model to make use of unlabeled data from the target domain, reducing the need for large labeled datasets and improving cross-domain performance.
Pure feature learning emphasizes directly extracting robust features from raw Wi-Fi data to improve performance across diverse domains without specific domain adaptation techniques. PAC-CSI (Su et al., 2023) develops a lightweight network to learn calibrated phase sequences for gesture recognition, incorporating digital twin techniques to simulate real-world scenes with 3D human models in Unity. 3D-ID (Ren and Yang, 2022) uses 2D AoA estimation to visualize individuals in physical environments, leveraging deep learning to create 3D body representations for person re-identification, using a Siamese network to ensure feature robustness. Abuhoureyah et al. (Abuhoureyah et al., 2023) proposes an adaptive algorithm that applies sequential learning to activity features, simplifying the recognition process and accommodating variations across different environments and individuals. Ma et al. (Ma et al., 2021) and (Sheng et al., 2020) integrate deep spatial-temporal models where LSTM captures temporal features and CNN extracts spatial features, providing a comprehensive understanding of static and dynamic patterns in Wi-Fi signals. These pure feature learning approaches improve Wi-Fi sensing models by focusing on extracting essential features, enabling robust performance across a wide range of tasks and domains.
2.4. Deployment Stage
Sensing models ultimately need to be deployed, and many studies focus on enhancing cross-environment generalization during deployment or enabling models to adapt to environmental changes and meet new task requirements through continual learning after deployment.
2.4.1. Transfer Learning with Fine-tuning
The typical steps of transfer learning involve first training a Wi-Fi sensing model on source domain data, followed by updating the model using target domain samples. This update can be applied to the task head, feature extractor, or the entire model. For example, XRF55 (Wang et al., 2024b), BullyDedect (Lan et al., 2024), and WiLISensing (Ding et al., 2020b) train models with source domain data and fine-tune the task head using several samples from the target domain, while freezing other parts of the model, enabling few-shot cross-domain activity recognition. DADA-AD (Chen et al., 2021) fine-tune the deep layers of the model. Besides, Brinke & Meratnia (Brinke and Meratnia, 2019b) experimentally compare the performance of fine-tuning the task head versus the entire network, finding that fine-tuning the whole network with 20% of target domain data yields the best results. In addition, in some cases, the target domain classification task differs from the source domain, such as in the number of action categories. To address this, WiTransfer (Fang et al., 2020) replaces the classification head and fine-tunes the new head using target domain data. FewSense (Yin et al., 2022), on the other hand, discards the classification head and fine-tunes the feature extractor using a cosine similarity loss. This loss pulls features of the same class closer while pushing features of different classes apart. During inference, it employs distance matching for classification, enabling the recognition of novel classes.
Another type of transfer learning consists of two stages: the meta-training stage and the meta-testing stage. In the meta-training stage, a rich labeled dataset from the source domain is used to train an embedding model or feature extractor. In the meta-testing stage, the embedding model is frozen, and a small number of samples from the target domain are input for feature extraction. These features are then used to fine-tune a task head, allowing the model to adapt to the target domain with limited data. Several typical works employ this framework, such as ResMon (Zheng et al., 2023), which uses BayesCNN as the embedding model for domain-adaptive respiration state monitoring. RoGER (Bi et al., 2024) leverages CNN as the feature extractor to achieve domain-robust gesture recognition. Hou et al. (Hou et al., 2022a) and DASECount (Hou et al., 2022b) both use CNN networks as feature extractors to achieve cross-domain crowd counting.
2.4.2. Federated Learning
WiFederated (Hernandez and Bulut, 2021), pFedBKD (Geng et al., 2023), CARING (Li et al., 2023), and Qi et al. (Qi et al., 2023) utilize federated learning frameworks to achieve efficient model deployment. Their core components include: (1) numerous clients with data collected from diverse environments, (2) a central server that trains a global model, which is expected to possess cross-environment generalization capabilities, and (3) the distribution of the global model to client devices, where it is fine-tuned using a small amount of local data. By leveraging these federated learning principles, Wi-Fi sensing models can be effectively deployed in unseen client environments.
Some works adopt frameworks similar to federated learning but use different terminology. For instance, Co-WiSensing (Zhang et al., 2023d) proposes a cloud-edge collaborative framework. In this approach, a cloud server trains a cloud Wi-Fi sensing model using data from various edge devices. The higher layers of the trained model are pruned, and the lower layers are distributed to the edge devices. At the edge, the pruned model is extended with a task-specific head, which is fine-tuned using the edge user’s local data to generate a Wi-Fi sensing model tailored for the edge environment. Zhang et al. (Zhang et al., 2021b) also employs a central server and adversarial learning to extract shared attributes across different environments. The activity-oriented attributes are then transferred to the local server, enabling cross-environment Wi-Fi human activity recognition. AdaWiFi (Zheng et al., 2024) proposes a collective sensing framework, where a master device collects data using multiple sensors and trains multiple encoders collaboratively, which are then aggregated into a unified model. In the end-user environment, the model is collaboratively tuned with a small amount of labeled data. Zhang et al. (Zhang et al., 2023f) propose a local–global modeling method for indoor localization. The approach involves clustering Wi-Fi fingerprints of a large environment into multiple sub-areas, training local localization models for each sub-area, and then aggregating these local models into a global model. The authors demonstrate that this method achieves superior cross-environment generalization.
2.4.3. Continuous Learning
After model deployment, it is desirable to continuously update the model to adapt to new environments or tasks. Some works achieve this with online learning (Shalev-Shwartz et al., 2012), a machine learning paradigm where the model learns and updates incrementally as new data becomes available. For instance, FewCS (Yang et al., 2023b) collects a small amount of online data to fine-tune model parameters in new environments, enabling few-shot cross-environment human action recognition. Similarly, WiDrive (Bai et al., 2019) utilizes online model adaptation and EM algorithm (Dempster et al., 1977) to update the model when in-car driver activities are misclassified, improving recognition for diverse vehicles and drivers. Unlike WiDrive’s automatic error detection, the User-in-the-loop approach allows users to report misclassifications or annotate a small number of samples, which are then used to update the model. M-WiFi (Soltanaghaei et al., 2020) specifically asks users to annotate critical time periods to fine-tune the trained model. To minimize the effort required from users, RISE (Zhai et al., 2021) combines probability and statistical assessments with anomaly detection to identify samples that are likely to be misclassified, reducing the need for users to report every error.
Class incremental learning and domain incremental learning are also applied to enable Wi-Fi human sensing models to adapt to new activities or environments. For example, CCS (Fu et al., 2024) uses a data replay-based incremental learning approach (Zhao et al., 2020), where the model is updated to sense new human activity categories using representative samples from previous stages and samples from the current stage. Similarly, CSI-ARIL (Zhang et al., 2023c) treats the environment as a domain, updating the model with representative samples from both old and current environments. Both CCS and CSI-ARIL ensure that the model does not suffer from catastrophic forgetting (Kirkpatrick et al., 2017) when adapting to new classes or domains, preserving its ability to sense old tasks or domains.
3. Dataset
Table 1. Open-sourced Dataset in Wi-Fi Human Sensing.
ID dataset (#E, #p, #A) Device setting Access
Action recognition
1 XRF55 (Wang et al., 2024b) 4 env., 39 part., 55 act. 1Tx, 3Rx, 20MB, 5GHz https://aiotgroup.github.io/XRF55/
2 Widar 3.0 (Zheng et al., 2019) 4 environment, 39 participants, 55 actions 1Tx, 3Rx, 20MB, 5GHz https://tns.thss.tsinghua.edu.cn/widar3.0/
3 ARIL (Wang et al., 2019a) 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/geekfeiw/ARIL
4 BullyDetect (Lan et al., 2024) 8 environments, 20 paired, 7 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/aiotgroup/WiFi-BullyDetect
5 HTHI (Alazrai et al., 2020) 1 environments, 40 paired, 12 actions 1Tx, 1Rx, 20MB, 5GHz https://data.mendeley.com/datasets/3dhn4xnjxw/1
6 WiAR (Guo et al., 2019) 5 environments, 10 participants, 16 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/linteresa/WiAR
7 Meneghello et.al. 2023  (Meneghello et al., 2023) 7 environments, 13 participants, 8 actions 1Tx, 1Rx, 20MB, 5GHz https://paperswithcode.com/paper/a-csi-dataset-for-wireless-human-sensing-on
8 Demrozi et.al. 2017 (Demrozi et al., 2023) 2 environments, 6 participants, 5 actions 1Tx, 1Rx, 20MB, 5GHz https://www.techrxiv.org/doi/full/10.36227/techrxiv.22284421.v1
9 Yousef et.al. 2017 (Yousefi et al., 2017) 1 environments, 6 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/ermongroup/Wifi_Activity_Recognition
10 WiSDA (Jiao et al., 2024) 1 environments, 5 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://www.researchgate.net/publication/385918549_WiSDA_Subdomain_Adaptation_Human_Activity_Recognition_Method_Using_Wi-Fi_Signals
11 WiNDR (Qin et al., 2024) 1 environments, 3 participants, 5 actions 3Tx, 3Rx, 20MB, 5GHz https://gitlab.com/yuxiqin/direction-independent
12 CSI-HAR-Dataset  (Moshiri et al., 2021) 3 environments, 6 participants, 5 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/parisafm/CSI-HAR-Dataset
13 WiMANS (Huang et al., 2024) 3 environments, 6 participants, 9 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/huangshk/WiMANS
14 ImgFi (Zhang and Jiao, 2023) 1 environments, 5 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://ieeexplore.ieee.org/document/10190332
15 ReWis (Bahadori et al., 2022) 3 environments, 2 participants, 3 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/niloobah/ReWiS
16 SAR (Brinke and Meratnia, 2019a, c) 1 environments, 9 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://data.4tu.nl/articles/_/12692816/1
17 DeepSeg (Xiao et al., 2020) 1 environments, 5 participants, 10 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/ChunjingXiao/DeepSeg
18 OPERAnet (Bocus et al., 2022) 2 environments, ??? participants, 7 actions 1Tx, 1Rx, 20MB, 5GHz https://springernature.figshare.com/collections/A_Comprehensive_Multimodal_Activity_Recognition_Dataset_Acquired_from_Radio_Frequency_and_Vision-Based_Sensors/5551209/1
19 NTU-Fi HAR (Yang et al., 2022a) 1 environments, 20 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/xyanchen/WiFi-CSI-Sensing-Benchmark
20 WiGuesture (Zhao et al., 2024b) 1 environments, 5 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/RS2002/CSI-BERT
21 RF-Net (Ding et al., 2020a) 6 environments, 11 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/di0002ya/RFNet
22 RISE (Zhai et al., 2021) 2 environments, 6 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/jiaojiao1234/RISE/
23 SignFi (Ma et al., 2018) 2 environments, 5 participants, 276 actions 3Tx, 1Rx, 20MB, 5GHz https://github.com/yongsen/SignFi
24 RoGER (Bi et al., 2024) 2 environments, 4 directions, 5 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz http://bisuzhi.cn/Sensedata.html
25 CSIDA (Zhang et al., 2021c) 2 environments, 5 locations, 5 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/Zhang-xie/WiGr
26 ResMon (Zheng et al., 2023) 2 environments, 3 areas, 6 participants, 4 actions 1Tx, 1Rx, 20MB, 5GHz http://bisuzhi.cn/Sensedata.html
27 Dataset3 (bis, 2024) 2 environments, 6 participants, 4 actions 1Tx, 1Rx, 20MB, 5GHz http://bisuzhi.cn/Sensedata.html
28 FallDeFi (Palipana et al., 2018) 5 environments, 3 participants, fall and others 1Tx, 1Rx, 20MB, 5GHz https://github.com/dmsp123/FallDeFi
29 FallDar (Yang et al., 2022b) 2 environments, 6 locations, 7 participants, fall and normal 1Tx, 1Rx, 20MB, 5GHz http://www.sdp8.net/Dataset?id=dfe4621e-4774-4ef4-b064-e4a082c12335
30 Baha’A et.al. 2020 (Baha’A et al., 2020) 3 environments, 50 participants, 5 actions 1Tx, 1Rx, 20MB, 2.4GHz http://www.sdp8.net/Dataset?id=dfe4621e-4774-4ef4-b064-e4a082c12335
31 SHARP (Meneghello et al., 2022) 3 environments, 4 locations, 3 participants, 4 actions 1Tx, 1Rx, 20MB, 2.4GHz https://github.com/francescamen/SHARP
32 CrossSense (Zhang et al., 2018) 3 environments, 15 locations, 100 participants, 40 actions 1Tx, 1Rx, 20MB, 5GHz Agreement requiredhttps://github.com/nwuzj/CrossSense/
Gait recognition
32 GaitID (Zhang et al., 2020c) 2 environment, 7 participants, 16 actions 20MB,5GHz http://www.sdp8.net/Dataset?id=87a65da2-18cb-4b8f-a1ec-c9696890172b
33 NTU-Fi HumanID (Wang et al., 2022e) 1 environment, 20 participants, 6 actions 1Tx, 3Rx,20MB,5GHz https://github.com/xyanchen/WiFi-CSI-Sensing-Benchmark
22 RISE (Zhai et al., 2021) 1 environments, 15 participants, walk 1Tx, 1Rx, 20MB, 5GHz https://github.com/jiaojiao1234/RISE/
30 CrossSense (Zhang et al., 2018) 3 environments, 100 participants, walk 1Tx, 1Rx, 20MB, 5GHz Aggrement required firsthttps://github.com/nwuzj/CrossSense/
Pose estimation
34 Person-in-WiFi 3D (Yan et al., 2024b) 3 environment, 7 participants, 8 actions 20MB,5GHz https://aiotgroup.github.io/Person-in-WiFi-3D/
35 $M$-Fi 4 environment, 40 participants, 27 actions 20MB,5GHz https://ntu-aiot-lab.github.io/$m$-fi
Indoor localization
3 ARIL (Wang et al., 2019a) 1 environment, 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz https://github.com/geekfeiw/ARIL
36 Chen and Chang 2022 (Chen and Chang, 2022) 2 environment, 7 participants, 16 actions 20MB,5GHz https://github.com/BingJiaChen/few-shot-indoor-localization
37 MetaLoc (Gao et al., 2023) 2 environment, 180 locations 20MB,5GHz https://github.com/StatFusion/MetaLoc
Crowd counting
38 DASECount (Hou et al., 2022b) 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MB, 5GHz http://bisuzhi.cn/Sensedata.html
Temporal Action Localization, Action Summarization
39 XRFv2 (Lan et al., 2025) 3 environment, 16 participants, 31 actions 1Tx, 3Rx, 20MB, 5GHz https://github.com/aiotgroup/XRFV2
Table 1. Open-sourced Dataset in Wi-Fi Human Sensing. (Continued.)
  1. (1)
    XRF55 (Wang et al., 2024b) dataset was collected in four indoor environments. In each environment, users performed 55 types of actions, categorized into five categories: human-object interactions, human-human interactions, fitness, body motions, and human-computer interactions, within a 3.1m $\times$× 3.1m area. The first scene involved 30 participants, while the other three scenes involved 3 participants each. Each participant performed each action 20 times, with 42,900 samples in total, lasting 59.58 hours. The data collection setup included one transmitter with a single antenna working, and three receivers each with three antennas, arranged in a quadrilateral layout. The sampling rate was 200 Hz. Notably, XRF55 dataset also includes synchronized modalities from millimeter-wave radar, RFID, and RGB+D+IR, supporting multimodal learning or cross-modal tasks.
  2. (2)
    Person-in-WiFi 3D (Yan et al., 2024b) dataset was collected in three environments: an office, a classroom, and a corridor, each within a 4m $\times$× 3.5m area. Seven participants freely performed eight actions: reaching out, raising hands, bending over, stretching, sitting down, lifting legs, standing, and walking. In addition to scenarios with a single participant, there were also cases where 2, 3, or 4 participants simultaneously performed actions within the scene. Each clip was recorded for 40 seconds, resulting in a total of 456 40-second recordings. The data collection setup included one transmitter with a single antenna and three receivers, each with three antennas, arranged in a quadrilateral layout. The sampling rate was 200 Hz.
  3. (3)
    ARIL (Wang et al., 2019a) dataset was collected in a 3.2m $\times$× 3.2m area within a laboratory, with 16 positions. The actions include 6 hand gestures for potential human-computer interaction applications: hand up, hand down, hand left, hand right, hand circle, and hand cross. One volunteer performed each gesture 15 times at each position. The dataset contains a total of 1,394 valid samples. Data was collected using two USRPx210 devices, each with a single antenna, as transmitter and receiver. The sampling rate was approximately 60Hz.
  4. (4)
    BullyDetect (Lan et al., 2024) dataset was collected in 8 different environments: a classroom, corridors, playgrounds, rooftops, a meeting room, a corner of the campus, an Office, and an abandoned building. Five volunteers were paired in 10 pairs, each acting as either an attacker or a victim during the experiment. The participants performed 7 different bullying actions: pushing, kicking, slapping, grabbing, punching, kneeing, and hitting with a stick. Each action was performed 5 seconds and 10 times, resulting in 11,200 samples. The experimental setup included a transmitter with a single antenna and one 3-antenna receiver, spaced 3 meters apart, with a sampling rate of 1000 Hz.
  5. (5)
[…]
  6. (6)
    GaitID (Zhang et al., 2020c) dataset is designed for gait recognition tasks. Data was collected in two scenarios: a hall and a discussion room. Volunteers walked along four predefined tracks, starting from both ends of each track. In the hall, 10 volunteers participated, with each walking 50 times in each direction. In the discussion room, 3 volunteers repeated each track 25 times in each direction. A total of 4,600 samples were collected. The data collection setup included one transmitter and six receivers. The sampling rate was not reporeted.
  7. (7)
    WiAR (Guo et al., 2019) dataset was collected in three indoor environments: one empty room, one meeting room, and one office, as well as two outdoor environments, including a playground and a corridor. Ten volunteers performed three categories of actions: upper body activities, lower body activities, and whole-body activities, with 16 different actions. Each action was performed 30 times. The experimental setup included a transmitter with a single antenna and one receiver with three antennas, spaced 4 meters apart. The participants performed the actions at the center of the setup. The sampling rate was 30Hz.
  8. (8)
    SignFi (Ma et al., 2018) dataset was collected in two environments: a laboratory and a home. The lab, measuring 13m $\times$× 12m, the distance between the AP and the STA was 230 centimeters. While the home environment, with dimensions of 4.11m $\times$× 3.86m, offered a shorter distance of 130 centimeters between the AP and STA. 5,520 instances were recorded in the lab, while the home environment provided 2,760 instances. The data involved 276 different sign language gestures, encompassing movements of the head, arm, hand, and fingers. The gestures were performed by 5 different users, and the users were instructed to perform each sign gesture repeatedly, with 20 instances for the lab environment and 10 instances for the home environment, resulting in a total of 8280 gesture instances.The sampling rate was 200Hz.
  9. (9)
    Meneghello et.al. 2023 (Meneghello et al., 2023) dataset was collected across seven diverse environments including a bedroom, living room, kitchen, university laboratory, university office, semi-anechoic chamber, and a meeting room. Thirteen participants consisting of ten males and three females freely performed eight actions: walking, running, jumping, sitting still, standing still, sitting downs, standing up, and arm exercises. And the dataset consists of more than 13 hours of data (amounting to 23.6 GB of raw CSI readings).
  10. (10)
[…]
  11. (11)
    Yousef et.al. 2017 (Yousefi et al., 2017)dataset was collected in an indoor office area where the WiFi transmitter and receiver were positioned 3m apart in a line-of-sight (LOS) condition. Six volunteers performed six different activities: lying down, falling down, walking, running, sitting down, and standing up. Each activity was performed 20 times, and each activity was recorded for 20 seconds. Additionally, video recordings were made during the experiments to help label and annotate the activities in the dataset.The sampling rate was 1 kHz.
  12. (12)
    WiSDA (Jiao et al., 2024) dataset was collected in an office setting, equipped with a desk and chair, within a $\times×2m$ sensing area. Five participants performed six activity categories: Push, Sweep, Tap, Slide, Draw-Z, and Draw-N. Each person collected 100 samples on 30 subcarriers and converted them to images, resulting in 3000 samples per action. The data was captured using a single pair of antennas, with one acting as the transmitter and the other as the receiver, positioned to face each other within the sensing area, similar to the setup used in the Widar 3.0 dataset.
  13. (13)
[…]
  14. (14)
    CSI-HAR-Dataset (Moshiri et al., 2021) dataset was collected in three environments incluing a classroom, a library, and a park. The dataset was collected by 6 volunteers including three male and three female. Each volunteer performed five categories of actions: sitting, standing, walking, running, and stretching. Sitting involved participants holding different postures on a chair for 5 seconds, while standing required them to maintain different body orientations for 5 seconds. Walking consisted of participants covering a distance of 10m at a normal pace, and running involved moving 10m at a moderate pace. Stretching included bending down, stretching arms overhead and so on. And each action was repeated 20 times by each participant. The experimental setup involved two cameras positioned 3m apart in the indoor spaces and 5m apart in the park. Additionally, a motion sensor was placed at the center of the test area, 4m away from the participants during indoor trials and 7m for the outdoor park trials. The sampling rate was 60Hz.
  15. (15)
    WiMANS (Huang et al., 2024) dataset was collected in three environments including a classroom, a meeting room and an empty room. Each environment was equipped with a transmitter, a receiver and a monitor camera used to capture synchronized videos. The activities were performed by five volunteers including an adult, a middle-aged person, and an elderly person. Nine activities were included in the dataset: nothing, walking, rotation, jumping, waving, lying down, picking up, sitting down and standing up. Each environment has five different locations. The dataset provides 11286 samples (over 9.4 hours) of dual-band WiFi CSI and synchronized videos.
  16. (16)
    ImgFi (Zhang and Jiao, 2023) dataset was collected in a laboratory, where 5 volunteers performed 6 actions: bending, drinking, nodding, squatting, drawing the letter O, and drawing the letter B. Each action was repeated 20 times, resulting in a total of 600 samples. Data collection was conducted using one transmitter and one receiver positioned 4 meters apart, with the volunteers performing actions in the center. The sampling rate was 100Hz.
  17. (17)
    ReWis (Bahadori et al., 2022) dataset was collected in three environments: an office, a meeting room, and a classroom. Two participants performed three actions—jumping, walking, and standing—10 times in each environment, with each session lasting 180 seconds and a minimum interval of 2 hours between sessions. A Netgear R7800 Wi-Fi router with a Qualcomm Atheros chipset operated in AP mode, while an off-the-shelf laptop functioned as the client. Additionally, three Asus RT-AC86U Wi-Fi routers, each equipped with four antennas, recorded Wi-Fi CSI at a rate of 100 Hz.
  18. (18)
    FallDeFi (Palipana et al., 2018), designed for fall detection, was collected in five environments: a kitchen, bedroom, corridor, laboratory, and bathroom. Three participants performed fall actions and other activities such as walking, jumping, sitting down, and standing up, with each action lasting for 10 seconds and 1153 samples in total. Each environment featured one transmitter and one receiver, with varying placement distances between 5 and 9 meters depending on the setting. To introduce environmental variations, data was collected on different days, and the transmitter’s position was adjusted during the experiments. All experiments were conducted at 5.2 GHz with a sampling rate of 1000 Hz.
  19. (19)
    FallDar (Yang et al., 2022b) dataset, developed for fall detection, was collected in two environments: a home environment and an office environment. In the home environment, five participants performed fall actions at three locations—dining room, living room, and balcony—resulting in 115 fall samples. The transmitter and receiver were placed in two settings: line-of-sight (LoS) and non-line-of-sight (NLoS). In the office environment, the transmitter was fixed in position, while the receiver was placed at 16 locations sequentially. Two participants contributed to the dataset. One performed fall actions at locations 1 and 3, while the other performed falls at locations 1 and 4. Additionally, one participant walked freely in the office to collect normal activity data. This resulted in 228 fall samples and 504 normal activity samples. In both environments, the transmitter had one antenna, and the receiver had three antennas. Data was collected at a sampling rate of 1,000 Hz with a center frequency of 5.825 GHz.
  20. (20)
    SAR  (Brinke and Meratnia, 2019a, c) dataset was collected in the living room of an apartment. A total of 9 volunteers participated in experiments over 6 days. During the first 3 days, 3 different volunteers participated each day, while 2 volunteers participated repeatedly over the last 3 days. During the first 3 days, participants performed 6 actions: Clap, Walk, Wave, Jump, Sit, and Fall. In the last 3 days, participants performed the same actions except for Jump, resulting in a total of 5 actions. Each action was performed continuously for 5 seconds and repeated 50 times. The data collection setup consisted of a transmitter and a receiver, each equipped with three antennas, spaced approximately 2.5 meters apart. The sampling rate was 50 Hz.
  21. (21)
    WiGuesture (Zhao et al., 2024b) dataset was collected in a conference room. Eight volunteers participated in the study. The participants performed six distinct gestures: moving left-right, forward-backward, up-down, circling clockwise, clapping, and waving. Each gesture was recorded continuously for one minute. The data collection setup consisted of a six-antenna home Wi-Fi router acting as the transmitter and a single-antenna ESP32S3 microcontroller serving as the receiver. The transmitter and receiver were placed 1.5 meters apart, and data was captured at a sampling rate of 100 Hz.
  22. (22)
    DeepSeg  (Xiao et al., 2020) dataset was collected in a meeting room with data from 5 volunteers performing 10 activities. These actions include 5 fine-grained movements such as swinging the hand, raising the hand, making a pushing motion, tracing a circular pattern (O), and tracing a cross pattern (X), as well as 5 coarse-grained activities, namely boxing, picking up an object, running, squatting down, and walking. Each participant repeated each action 30 times, resulting in 1,500 samples. The dataset collection setup included a single-antenna transmitter and a three-antenna receiver placed 2 meters apart, with a sampling rate of 50 Hz.
  23. (23)
    OPERAnet (Bocus et al., 2022) dataset was collected in two rooms and includes data for seven activities: walk, sit, stand from a chair, lie down, stand up from the floor, rotate the upper-half body, and steady state, totaling 417 minutes of activity data. Additionally, the dataset includes 18 minutes of background data and 27 minutes of crowd-counting data. The data collection setup consisted of one transmitter with three antennas and two receivers, each with three antennas, arranged in a right-angled triangular configuration. The second receiver was positioned at the right-angle vertex of the triangle. The sampling rate for Wi-Fi signals was 1600 Hz. OPERAnet dataset also includes synchronized data from two Kinect cameras, four UWB nodes, and a Passive WiFi Radar system constructed using a USRP-2945. This multimodal setup supports research in both multi-modal learning and cross-modal tasks.
  24. (24)
    Chen and Chang (Chen and Chang, 2022) dataset is used for indoor positioning based on CSI (Channel State Information) fingerprints and includes two environments: conference and office. In the conference scene, there is one pair of transmitters and receivers, and data is collected at 16 different locations, with 600 CSI fingerprint samples recorded at each location, resulting in a total of 9600 samples. In the office scene, two pairs of transmitters and receivers are used, with data collected at 18 locations, and 500 CSI fingerprint samples recorded at each location, yielding a total of 9000 samples.
  25. (25)
    HTHI (Alazrai et al., 2020) dataset was collected in a furnished room. 66 volunteers participated in the study, forming 40 unique pairs to perform 12 types of human-to-human interaction activities, such as handshaking, high fives, and hugging. Each pair performed each activity 10 times, resulting in 4,800 samples. The data collection setup consisted of a commercial off-the-shelf Sagemcom 2704 router with two antennas acting as the transmitter and a desktop computer with three antennas serving as the receiver. The transmitter and receiver were positioned 4.3 meters apart. The sampling rate for the data collection was not reported.
  26. (26)
    MM-Fi (Yang et al., 2024a) dataset comprises data collected in two rooms, 4 environmental settings, with 10 participants in each environment. The participants performed 14 daily activities, such as raising the left arm, waving the left arm, and waving the right arm, as well as 13 rehabilitation exercises, such as left front lunge and right front lunge. The setup included a transmitter with one antenna and a receiver with three antennas, arranged with a spacing of 3.75 meters between the receiver antennas. Participants performed actions while positioned 3 meters away from the receiver, facing the receiver. Each action was performed for 30 seconds. The devices operated at 5 GHz with a sampling rate of 1000 Hz. In addition to Wi-Fi sensing, MM-Fi includes data from LiDAR, mmWave radar, and camera modalities, supporting multimodal learning tasks.
  27. (27)
[…]
  28. (28)
    NTU-Fi HAR (Yang et al., 2022a) was collected in a laboratory setting with 20 participants performing six types of actions, including running, walking, falling down, boxing, circling arms, and cleaning the floor. Each participant performed each action 20 times, resulting in a total of 2,400 samples. The setup consisted of a transmitter with one antenna and a receiver with three antennas. The equipment operated at 5 GHz with a sampling rate of 500 Hz.
  29. (29)
    CrossSense (Zhang et al., 2018) dataset was collected in three scenarios: a hall entrance, a narrow corridor, and a room, with the participation of 100 volunteers. For gait recognition, each volunteer walked 20 times in the same direction in each scenario, resulting in a total of 6,000 samples. For action recognition, five predefined positions were set in each scenario. At each position, each volunteer performed 40 different actions (e.g., pull, drag, throw), repeating each action 10 times. This resulted in a total of 600,000 samples. The experimental setup included one transmitter and one receiver, both equipped with three antennas. The devices operated at 5 GHz with a sampling rate of 1,000 Hz. Currently, access to the CrossSense dataset requires completing a consent form in its github repository.
  30. (30)
    RF-Net (Ding et al., 2020a) dataset was collected in six different rooms with 11 volunteers performing six types of actions: wiping, walking, moving, rotating, sitting, and standing up. The Wi-Fi setup included two transmitter-receiver pairs, operating at a sampling rate of 100 Hz.
  31. (31)
    CSIDA, related in WiGr (Zhang et al., 2021c), was collected in an office environment with three distinct locations and a classroom environment with two locations. Five volunteers performed six gestures: pull left, pull right, lift up, press down, draw a circle, and draw zigzag. Each gesture was repeated 10 times, with each repetition lasting 1.8 seconds. The device setup consisted of a single-antenna transmitter and a three-antenna receiver, separated by a distance of 2.6 meters. The system operated in the 5 GHz mode with a sampling rate of 1000 Hz.
  32. (32)
    RISE (Zhai et al., 2021) dataset includes data for both gesture recognition and gait recognition, collected in two environments: an office and a controlled setting featuring an RF anechoic chamber to minimize multipath effects. Gesture Recognition: six volunteers performed six gestures—push and pull, draw a circle, throw, slide, sweep, and draw zigzag—across five locations in both environments. Each participant repeated each gesture 30 times at each location, resulting in a total of 10,800 samples. Gait Recognition: 15 volunteers walked under five different configurations in the controlled environment. Each volunteer completed 30 walks, generating 1,150 samples in total. The setup consisted of one transmitter and one receiver, operating at a sampling rate of 1,000 Hz.
  33. (33)
    RoGER (Bi et al., 2024) dataset was collected in a meeting room at two distinct locations. Location 1: A single volunteer performed six gestures, i.e., Push & Pull, Draw Zigzag, Clap, Sweep, Draw Circle, Slide, the same as Widar 3.0 (Zheng et al., 2019), in four different directions. Each gesture was repeated 30 times, resulting in 480 samples. Location 2: Four volunteers (including the participant from Location 1) performed the same six gestures in four directions. Each volunteer repeated each gesture 30 times, generating 1,920 samples. Each sample lasts for 10 seconds. The setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.437 GHz with a sampling rate of 100 Hz.
  34. (34)
    ResMon (Zheng et al., 2023) dataset was collected in two environments: a laboratory and a meeting room, with the laboratory further collected at two areas. Six volunteers performed 4 actions: stable breath, cough, sneeze, and yawn. The experimental setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.437 GHz with a sampling rate of 100 Hz.
  35. (35)
    DASECount  (Hou et al., 2022b) dataset, designed for crowd counting, was collected in two environments: an office and a lecture hall, each with both LOS and NLOS configurations, resulting four different environmental settings. Volunteers, ranging from 0 to 8 people, performed three types of motion: static, dynamic, and mixed. Each sampling sequence was recorded for 5 minutes. The experimental setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.437 GHz with a sampling rate of 100 Hz.
  36. (36)
    Dataset3 (bis, 2024) is used for fatigue detection and includes two environments: a meeting room and an office room. It contains four types of fatigue actions: yawning, rubbing eyes, nodding, and stillness. Six volunteers performed each action 50 times in the meeting room and 40 times in the office room, resulting in a total of 2160 samples from both environments.
  37. (37)
    MetaLoc (Gao et al., 2023) dataset was designed for indoor localization and collected in a 12 m $\times$× 5 m hall and a 10 m $\times$× 8 m laboratory. In each environment, Wi-Fi fingerprints were recorded at 90 locations. The setup included three transmitters: an ASUS RT-AC86U, a TP-Link TL-WR885N, and a TP-Link TL-WR886N, and a Nexus-5 smartphone as the receiver. The system operated at 5 GHz.
  38. (38)
    Baha’A et.al. 2020 (Baha’A et al., 2020) data collected in three different environments: two with line-of-sight (LOS) settings and one with a non-line-of-sight (NLOS) setting. Thirty volunteers performed five different actions: falling from a sitting position, falling from a standing position, walking, sitting down and standing up, and picking up a pen from the ground. Each action was performed 20 times. The experimental setup included a single-antenna transmitter and a three-antenna receiver, operating at a frequency of 2.4 GHz with a sampling rate of 320Hz.
  39. (39)
    SHARP (Meneghello et al., 2022) dataset was collected in three environments: a bedroom, a living room, and a laboratory, featuring four activities—sit, walk, run, and jump—along with an empty room state. In the bedroom, volunteers p1 and p2 performed at two locations; in the living room, volunteer p1 performed at one location; and in the laboratory, volunteer p3 performed. These activities were conducted at different times over ten months. Each activity lasted 120 seconds per session, resulting in a total of 120 minutes of data. The setup included a single-antenna transmitter and receiver, along with a 4-antenna Asus RT-AC86U router as a monitoring device for collecting Wi-Fi data at a sampling rate of 173 Hz.
  40. (40)
    XRFv2 (Lan et al., 2025) dataset is designed for continuous action localization and action summarization tasks. It was collected in three distinct indoor settings—a study room, a dining room, and a bedroom. In each environment, 16 volunteers performed continuous sequences of actions spanning 30 categories. The dataset comprises 853 continuous action sequences, each lasting approximately 80 seconds, making it suitable for continuous action detection, localization, and summarization tasks. The data collection setup utilized one transmitter and three receivers with a sampling rate of 200 Hz. In addition to Wi-Fi measurements, participants wore six IMU sensors placed on an earphone, glasses, two phone pockets, and two wrist-worn watches. Furthermore, XRFv2 provides pose annotations, supporting multimodal fusion and cross-modal research.
4. Challenges and Future Directions
4.1. Data, Data, Data
The scaling law, which involves training large-scale models with large-scale datasets, has proven its success since the advent of AlexNet (Krizhevsky et al., 2012), achieving remarkable breakthroughs in fields such as computer vision (CV), natural language processing (NLP), robotics, and science. Large-scale datasets like ImageNet (Deng et al., 2009), COCO (Lin et al., 2014), Kinetics (Kay et al., 2017), and KITTI (Geiger et al., 2012) have significantly accelerated advancements in these areas. In contrast, the field of Wi-Fi human sensing has not progressed as rapidly, despite the widespread adoption of deep learning methods. A key reason lies in the limited scale and diversity of available Wi-Fi sensing datasets.
Although researchers have made significant efforts to close this gap by collecting and releasing datasets like XRF55 (Wang et al., 2024b) and Widar 3.0 (Zheng et al., 2019), these datasets still suffer from constraints in terms of collection scenarios, action categories, and participant diversity. When compared to the scale of video, image, or text datasets, Wi-Fi sensing datasets lag behind by several orders of magnitude.
Collecting and annotating Wi-Fi sensing data is inherently time-consuming and labor-intensive. To address this bottleneck, synthetic data generation emerges as a promising solution. By producing large volumes of realistic data, synthetic methods can support the training of models with substantial parameters, potentially revolutionizing Wi-Fi sensing research. However, several open challenges remain to be addressed:
  1. (1)
    How to model the propagation of signals through people and environments? Can methods like ray-tracing and Fresnel zone models effectively simulate Wi-Fi signals’ interactions with objects and people? How to enable modern approaches, such as diffusion models (Ho et al., 2020), Variational Autoencoders (VAEs) (Kingma and Welling, 2013), Neural Radiance Fields (NeRF) (Mildenhall et al., 2021), or 3D Gaussian Splatting (Kerbl et al., 2023) learn probabilistic mappings of signal propagation patterns from observed data?
  2. (2)
    How to construct diverse environments for data generation? Should we rely on manual modeling of environments, even though it is resource-intensive? Or should we leverage open-source tools like Unity, Unreal Engine, or Infinigen Indoors (Raistrick et al., 2024), which provide scalable solutions for generating dynamic, realistic environments tailored to Wi-Fi sensing scenarios?
  3. (3)
    How to synthesize diverse human body models? Is it more effective to use game engines like Unity or Unreal Engine to create diverse human models? Alternatively, could tools such as SMPL (Loper et al., 2015) or MANO (Romero et al., 2022) be utilized to extract human body models from large-scale video datasets for synthetic data generation?
  4. (4)
    How to bridge the gap between synthetic and real data? Could Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) help enhance the realism of synthetic features or adapt synthetic data to better match real-world characteristics? What role could domain adaptation and domain-invariant feature learning play in aligning synthetic and real data to improve model performance?
Addressing these challenges offers a pathway to creating realistic, diverse, and scalable synthetic Wi-Fi sensing datasets. By bridging the data gap, the Wi-Fi sensing field could achieve progress comparable to the rapid advancements seen in video, image, and text recognition, ultimately unlocking new frontiers for this promising domain.
4.2. Wi-Fi Sensing Foundation Model Pre-Training
Foundation models in vision and language, such as CLIP (Radford et al., 2021), GPT (Brown et al., 2020), and BERT (Devlin et al., 2019), have demonstrated remarkable success in adapting to diverse downstream tasks. Similarly, Wi-Fi sensing encompasses a wide range of applications, including human activity recognition, indoor localization, presence detection, and pose estimation. These varied tasks make the development of a Wi-Fi sensing foundation model both promising and necessary. However, training such a model raises several critical open questions that must be addressed:
  1. (1)
    Should we utilize traditional architectures like CNNs and RNNs, or opt for Transformers (Waswani et al., 2017), which have demonstrated strong performance across various domains? Alternatively, emerging architectures such as Mamba (Gu and Dao, 2023), known for their smaller parameter sizes, reduced memory consumption, and faster inference speeds, may also be viable options. Should we further tailor the architecture design to cater to different computational platforms, optimizing for specific hardware constraints and efficiency requirements?
  2. (2)
    How can unlabeled real-world data be incorporated effectively? While large-scale synthetic labeled data and small amounts of labeled real-world data may be used for training, how can we leverage large-scale unlabeled real-world data in the pre-training of the foundation model? Should this real-world data be transmitted to a central server for processing, or should edge computing be adopted to process data locally? Additionally, how can we coordinate with a large number of real-world users to ensure efficient data collection and processing while respecting privacy and computational constraints?
  3. (3)
    What self-supervised learning strategy is most suitable? Self-supervised learning can effectively leverage unlabeled data, but how should we design the proxy tasks? Should we employ contrastive learning strategies that bring positive sample pairs closer while pushing negative pairs apart? Or would a masking-then-reconstruction approach, as used in models like BERT (Devlin et al., 2019) or MAE (He et al., 2022), be more appropriate for Wi-Fi sensing? Additionally, are there domain-specific proxy tasks that could better capture the unique spatiotemporal characteristics of Wi-Fi signals?
  4. (4)
    Wi-Fi sensing is significantly influenced by environmental factors and device configurations. Incorporating such information could enable the foundation model to reason more effectively about environmental changes and adapt to new scenarios. Should this integration be achieved through explicit embeddings of environmental and device parameters, or should the model be designed to infer these characteristics implicitly from the data?
Addressing these questions will be critical for developing a robust and generalizable Wi-Fi sensing foundation model that can serve as a backbone for various downstream applications.
4.3. When Wi-Fi Sensing Meets Large Multimodal Models
Beyond training Wi-Fi foundation models from scratch, as discussed in Section 4.2, another promising research direction is leveraging existing large multimodal models by fine-tuning them to process Wi-Fi sensing data. Large multimodal models, such as those trained on diverse combinations of images, text, and audio, have demonstrated impressive adaptability across various domains. By aligning Wi-Fi sensing data with these models, researchers may unlock new possibilities for enhanced understanding and application of Wi-Fi sensing technologies. However, integrating Wi-Fi data into such models raises several open questions and challenges:
  1. (1)
    What fine-tuning strategy is most effective? Fine-tuning large multimodal models requires balancing computational efficiency with task-specific performance improvements. Techniques like LoRA (Hu et al., 2022), adapters (Rebuffi et al., 2017), or prompt-based fine-tuning (Liu et al., 2021) are popular in the NLP and vision domains. Should one of these strategies be adopted for Wi-Fi sensing, or is there a need for a novel approach tailored to the unique characteristics of Wi-Fi data? Moreover, how can we efficiently fine-tune large models with limited computational resources while ensuring high adaptability?
  2. (2)
    How can Wi-Fi data be aligned with the modalities handled by large multimodal models? Wi-Fi sensing data typically exists in time-series or signal-based formats, which differ significantly from image or text data commonly processed by multimodal models. Should the raw Wi-Fi data be transformed into spectrograms or other image-like representations for easier integration, or is it better to retain the time-series format and develop alignment techniques specific to such data? Furthermore, what pre-processing or feature extraction methods are necessary to achieve optimal performance when aligning Wi-Fi data with existing modalities?
  3. (3)
    How should domain-specific knowledge be incorporated into the model? Wi-Fi sensing encompasses a wealth of domain-specific knowledge, including signal propagation principles, communication protocols, e.g. IEEE 802.11 bf, and established research findings. Should this information be used to fine-tune models before fine-tuning on Wi-Fi signals, or would integrating Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) as an external database allow the model to access Wi-Fi knowledge on demand? How can this domain knowledge improve interpretability, efficiency, and overall performance?
  4. (4)
    What role does multimodal fusion play in Wi-Fi sensing? Large multimodal models excel at integrating and reasoning over diverse data sources. In Wi-Fi sensing, how can complementary modalities such as video, audio, or inertial measurements (e.g., IMU data) be effectively fused with Wi-Fi data? What techniques can ensure seamless fusion while addressing challenges like modality-specific noise, synchronization issues, and computational overhead?
Addressing these challenges will advance the integration of Wi-Fi sensing with large multimodal models and push the boundaries of their application across various tasks and environments. This direction holds great potential for accelerating progress in Wi-Fi sensing by leveraging the capabilities of state-of-the-art AI models.
4.4. System Deployment and Service Models
As Wi-Fi sensing technologies advance and large multimodal models are integrated into Wi-Fi sensing applications, an essential next step involves addressing deployment and service models for these systems. Several critical questions arise when considering how to effectively and efficiently deploy such systems in real-world settings:
  1. (1)
    Large multimodal models often require high computational power and storage. Compression techniques like pruning (Han et al., 2015), quantization (Jacob et al., 2018), or knowledge distillation (Hinton et al., 2015) can help balance performance and efficiency. How can these methods be tailored for different hardware (e.g., edge devices or mobile phones) to ensure speed without sacrificing too much performance?
  2. (2)
    How to manage network bandwidth, data traffic, and latency during operation? Network constraints, particularly with real-time data transmission, pose challenges. How can continuous data flow between edge devices and servers be optimized? Strategies like edge computing (Shi et al., 2016), which processes data locally, could reduce latency and bandwidth usage but introduce complexity in resource management and synchronization.
  3. (3)
    Should model updates be centralized or on-demand at the edge? Model updates can either be pushed centrally from a server or performed on-demand at the edge based on local needs. Centralized updates ensure consistency but may face bandwidth and computation constraints. Edge updates offer quick adaptability but can complicate coordination, version control, and synchronization.
  4. (4)
    What service and pricing strategies should be adopted? Deciding on the right service and pricing models is vital for commercialization. Should services be cloud-based, subscription-based, or pay-per-use? How can pricing be adapted based on usage, such as the number of devices, frequency of access, or task complexity? Should tiered services be offered, with basic models for simpler tasks and premium ones for more complex applications?
Addressing these questions will help define how Wi-Fi sensing services can be deployed and scaled efficiently in real-world applications, ultimately determining the practicality and success of Wi-Fi sensing systems across diverse industries and environments.
4.5. Security and Privacy
The integration of Wi-Fi sensing in various applications introduces several security and privacy concerns. These issues need to be addressed to ensure the safe and ethical use of Wi-Fi sensing technologies in real-world scenarios. Some of the key challenges include:
  1. (1)
    How can sensitive user data be handled effectively? Wi-Fi sensing data may inadvertently capture sensitive information, such as user identity or pose (Wang et al., 2019b, f). How can we ensure that such details are not exposed in applications like crowd counting, where they are unnecessary? What strategies can be employed to anonymize or remove sensitive data without compromising the performance of task-specific models? Can techniques such as differential privacy, data anonymization, or task-specific filtering be used to ensure that only relevant information is retained while protecting user privacy?
  2. (2)
    How can sensing models be made resilient to attacks? Sensing models may be vulnerable to adversarial attacks that disrupt or manipulate the system output (Liu et al., 2022). How can we protect Wi-Fi sensing systems from attackers who try to interfere with the communication channel and alter the sensing results? What approaches, such as robust model training, error correction, and secure data transmission protocols, can be adopted to mitigate such risks and ensure the reliability of applications?
  3. (3)
    What privacy concerns arise from Wi-Fi penetration ability? Wi-Fi sensing has the potential to see through walls (Adib and Katabi, 2013) and capture the behavior of individuals in adjacent rooms or neighboring environments. How can we address privacy leakage concerns, particularly in residential or private spaces? What privacy-preserving protocols need to be designed to limit the scope of sensing to the intended areas? Should regulations or hardware modifications be implemented to restrict the range of Wi-Fi signals and ensure that sensitive information is not accessed without consent?
In conclusion, as Wi-Fi sensing technologies advance, it is imperative to integrate robust security and privacy measures that protect users and ensure the ethical use of the technology. Developing secure models, using privacy-preserving techniques, and complying with relevant regulations will be key to the success and wide adoption of Wi-Fi sensing applications.
5. Conclusion
This survey encompasses a significantly broader range of Wi-Fi generalization studies—over 200 papers in total. We have organized these studies along the entire Wi-Fi sensing pipeline: from device deployment and signal processing to feature learning and model deployment. This structured approach enables researchers to quickly locate the techniques that best address their specific challenges. Furthermore, we provide a detailed overview of open-source datasets, helping researchers efficiently identify the data needed to validate their algorithms. Finally, we offer a comprehensive analysis of future directions, including trends in multimodal approaches and large language models, to inspire new research ideas. In summary, this work serves as a valuable resource for researchers seeking convenient access to methodologies, datasets, and innovative research strategies.
Contributions
Fei Wang initiated and led the project, reviewed and organized all the papers on Wi-Fi generalizability and datasets, and was responsible for writing the manuscript. Tingting Zhang reviewed and organized approximately one-fifth of the generalizability and dataset papers, participating in discussions on the manuscript. Bintong Zhao, Libao Xing, and Tiantian Wang each reviewed and categorized approximately one-fifth of the Wi-Fi generalizability papers.